OpenStudy (mollykenney):

how do i find the slope intercept form for the points (-8,6) (1,-9)

jimthompson5910 (jim_thompson5910):

first find the slope of the line through the two points m = (y2-y1)/(x2-x1) m = (-9-6)/(1-(-8)) m = (-9-6)/(1+8) m = (-15)/9 m = -5/3

jimthompson5910 (jim_thompson5910):

now turn to y = mx+b form, plug in the slope and one given point, then solve for y = mx+b y = (-5/3)x + b ... plug in the slope 6 = (-5/3)(-8) + b ... plug in the first point 6 = 40/3 + b solve for b, I'll let you finish up

OpenStudy (mollykenney):

so to solve for b i will just bring the 40/3 and end up subtracting 6- 40/3 then i will be done?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

6 - 40/3 is what

OpenStudy (mollykenney):

-22/3

jimthompson5910 (jim_thompson5910):

perfect

jimthompson5910 (jim_thompson5910):

so y = mx + b turns into \[\large y = -\frac{5}{3} - \frac{22}{3}\]

OpenStudy (mollykenney):

but you still leave space to plug in x right?

OpenStudy (mollykenney):

such as y= -5/3 (x) - 22/3

jimthompson5910 (jim_thompson5910):

oh sry, meant to say \[\large y = -\frac{5}{3}x - \frac{22}{3}\]

jimthompson5910 (jim_thompson5910):

yeah you got it

OpenStudy (mollykenney):

perfect thanks so much

jimthompson5910 (jim_thompson5910):

you're welcome